OpenStudy (anonymous):

I just need someone to explain how to do this please. I also have a few more questions that go along with it if you're willing to help with those too please. In order to earn an A in class, your test scores must average at least 92%. On the first 6 tests, you have an average of 91%. What score do you need to get on the last test to earn an a?

OpenStudy (anonymous):

6*91=546 546+98=644 644/7=92

OpenStudy (anonymous):

So, you need a 98 on the last test to get a 92

OpenStudy (anonymous):

Recall what the average really is, to calculate it, you add up all the numbers in the set and divide by how many numbers you added up, alright, that means whatever scores you got on the first 6 tests, they result in this\[\frac{ x_1+x_2+x_3+x_4+x_5+x_6 }{ 6 }=.91\]for simplicity, lets just say they scores .91 on all the previous 6 tests, well now you want to add in another score, so we are caculating a new mean that looks like this,\[\frac{ x_1+x_2+x_3+x_4+x_5+x_6+x_7 }{ 7 }=.92\]but we know that the first 6 scores can be written as\[6(.91)\]that would make the equation look more like this\[\frac{ 6(.91)+x }{ 7 }=.92\]anyone can correct me if i'm making any mistakes, but this seems logical.

OpenStudy (anonymous):

At least a 98% on the last test to get an A.

OpenStudy (anonymous):

Looks right to me. Very explanatory.

OpenStudy (anonymous):

Thank God for people who are good at math. Sadly I wasn't blessed with this trait. Would any of you guys help me with about 3 more?? I promise you're not doing my entire homework for me like half the other people on here. It's a 7 page packet and these are the only ones I can't seem to figure out to save my life.